OpenStudy (anonymous):

[4_01]Which of the following is the slope between the points (0, -10) and (-4, 0)? 2/5 -2/5 -5/2 5/2

OpenStudy (anonymous):

slope point formula 0-0/-4-(-10) = slope 0/14= 0 slope is a line y=o

OpenStudy (anonymous):

sorry where did you get choices ....

OpenStudy (anonymous):

those are my options to choose from

OpenStudy (isaiah.feynman):

0-(-10)/-4-0 = 10/-4= -5/2

OpenStudy (isaiah.feynman):

So -5/2

OpenStudy (anonymous):

read it wrong 0-(-10)/ -4-0 = 10/4 =5/2

OpenStudy (anonymous):

Yep it is negative

OpenStudy (anonymous):

Thanks :-) so its -5/2 ?

OpenStudy (anonymous):

you are done with this question

OpenStudy (isaiah.feynman):

@aleska_t its -5/2